# Uplink-Downlink Duality for Beamforming in Integrated Sensing and Communications

### v1

@@ III-B Analysis of the Dual Problem
The reformulated problem (30) is convex with strong duality. We proceed by taking its Lagrangian dual with respect to (29b). Let $\tilde{\mathbf{B}}_{1},\ldots,\tilde{\mathbf{B}}_{L}\in\mathbb{R}^{(L+1)\times(L+1)}$ denote the dual variables associated with (29b)
@@
which holds because $\mathbf{J}_{\mathbf{R}}$ is PSD. Thus, the optimal $\mathbf{B}_{\ell}$ must be $\mathbf{B}_{\ell}=\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}^{\mathsf{T}}_{\ell}$ at the saddle-point. Substituting this relation into (33), and again interchanging maximization and minimization, we arrive at
@@
Let $\mathcal{V}$ denote the constraints (12b)-(12c). Suppose the optimization problem (12) is strictly feasible and that Assumption 1 holds. Then, $\mathbf{V}^{*}$ is a global solution of (12) if and only if there exists $\boldsymbol{\beta}^{*}\in\mathbb{R}^{L\times L}$ such that $(\mathbf{V}^{*},\boldsymbol{\beta}^{*})$ is a saddle-point solution of
@@
where $\boldsymbol{\beta}\triangleq[\boldsymbol{\beta}_{1},\ldots,\boldsymbol{\beta}_{L}]\in\mathbb{R}^{L\times L}$.
@@
$\mathbf{Q}_{\boldsymbol{\beta}}\triangleq\frac{2T}{\sigma^{2}}\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right],$ (40)
@@
The development leading to the theorem shows that (12) is equivalent to (38). To further expand the term involving $\mathbf{J}_{\mathbf{V}}$, we note from (6) that
@@
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\sum_{ij}[\boldsymbol{\beta}_{\ell}]_{i}\mathbb{E}_{\boldsymbol{\eta}}\left[({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}[\boldsymbol{\beta}_{\ell}]_{j}\right)$ (43b)
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[\sum_{i}[\boldsymbol{\beta}_{\ell}]_{i}({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\sum_{j}[\boldsymbol{\beta}_{\ell}]_{j}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43c)
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43d)
@@
Examining the max-min problem (39), by construction the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$ is PSD for any $\boldsymbol{\beta}$. Consequently, the trace term can be interpreted as a sum of beamforming powers measured with respect to the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$. Thus, for a fixed $\boldsymbol{\beta}$, the inner problem in (39) is a power maximization along certain directions of interest. Varying $\boldsymbol{\beta}$ corresponds to varying the directions of interest (as illustrated in Fig. 2). Theorem 1 states that there exists a judicious choice $\boldsymbol{\beta}^{*}$ for which the problem of minimizing the BCRB is equivalent to that of maximizing the power in the direction of $\mathbf{Q}_{\boldsymbol{\beta}^{*}}$.
@@ III-C AoA Estimation Example
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}=\frac{2T}{\sigma^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (45)
@@ IV-A ISAC as Weighted Downlink Power Minimization
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (49b)
@@
$\mathrm{s.t.\ }$~~~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k.$ (51b)
@@
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k.$ (52b)
@@
This PSD condition on $\mathbf{\tilde{Q}}$ is important because the main result of [12] is that the weight matrix $\mathbf{\tilde{Q}}$ for the downlink power minimization problem becomes the noise covariance matrix in the dual uplink channel. Without the condition $\mathbf{\tilde{Q}}\succcurlyeq 0$, it cannot be a valid noise covariance.
There is no guarantee that for arbitrary $\lambda$ and $\boldsymbol{\beta}$, we would have $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$. Instead, we define the following notion of admissibility to restrict the space of $(\lambda,\boldsymbol{\beta})$ under consideration. It turns out that the optimal $(\lambda^{*},\boldsymbol{\beta}^{*})$ is always admissible, and further, an uplink-downlink duality relationship can be established for all admissible $(\lambda,\boldsymbol{\beta})$’s.
@@
The pair $(\lambda,\boldsymbol{\beta})$ is said to be admissible for a set of channels $(\mathbf{h}_{1},\cdots,\mathbf{h}_{K})$ under the SINR constraints $(\gamma_{1},\cdots,\gamma_{K})$ for the $K$ users with noise variance $\sigma^{2}$, if whenever a set of $\mathbf{V}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}]$ satisfies
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (53)
@@
Admissibility of $(\lambda,\boldsymbol{\beta})$ implies that the matrix $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ behaves somewhat like a PSD matrix, but the condition is much weaker than PSD, not only because $\mathbf{v}_{k}$’s are restricted to the feasibility set satisfying the SINR constraints, but also only the sum power is required to be nonnegative. The individual term $\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}$ can actually be negative.
@@
More importantly, we can show that whenever $(\lambda,\boldsymbol{\beta})$ is admissible, an uplink-downlink duality result can be established for the inner minimization in (51), or equivalently (52), as explored in detail in the next section. Such uplink-downlink duality would have been a direct consequence of prior work [12] if $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$, in which case $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ becomes the noise covariance in the dual uplink channel. The interesting and nontrivial fact is that even if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not PSD (so the uplink noise variance may be negative) as long as $(\lambda,\boldsymbol{\beta})$ is admissible, a duality relationship can still be established.
@@ IV-B Uplink-Downlink Duality
Let $\mathcal{A}$ be the set of admissible $(\lambda,\boldsymbol{\beta})$. The following theorem states an uplink-downlink duality relationship under the condition that $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. The dual uplink channel here has to satisfy an additional technical condition related to a certain matrix being an $M$-matrix.
@@
Consider the downlink beamforming problem for fixed $(\lambda,\boldsymbol{\beta})$
@@
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (55b)
where the constraints are assumed to be strictly feasible. The optimal value of this downlink optimization problem is bounded below if and only if $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$.
Furthermore, for $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, an optimal solution $\mathbf{V}^{*}_{\lambda,\boldsymbol{\beta}}\triangleq\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ to (55) can be obtained as
$\mathbf{V}^{*}_{\lambda,\boldsymbol{\beta}}=\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}\mathbf{p}_{\lambda,\boldsymbol{\beta}}^{*},$ (56)
where $\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}\in\mathbb{C}^{N_{\text{T}}\times K}$ is a matrix of uplink beamformers obtained by solving the following dual uplink problem over the normalized uplink receive beamformers $\mathbf{U}\triangleq\left[\mathbf{u}_{1},\ldots,\mathbf{u}_{K}\right]$ and the uplink transmit powers $\mathbf{q}\triangleq\left[q_{1},\ldots,q_{K}\right]^{\mathsf{T}}\in\mathbb{R}^{K}$:
$\underset{\mathbf{U},\;\mathbf{q}}{\mathrm{minimize}}$~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (57a)
$\mathrm{subject\ to}$~ $\frac{q_{k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}}{\sum_{i\neq k}q_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}+\mathbf{u}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{u}_{k}}\geq\gamma_{k},\forall k$ (57b)
@@
$\mathbf{M}_{\mathbf{U}}^{-1}\geq 0,$ (57d)
and $\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}\in\mathbb{R}^{K}$ is a vector of downlink powers obtained from
$\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}=\sigma^{2}\mathbf{M}_{\mathbf{U}_{\lambda,\boldsymbol{\beta}}^{*}}^{-1}\mathbf{1},$ (58)
where $\mathbf{M}_{\mathbf{U}}\in\mathbb{R}^{K\times K}$ is a matrix defined as
@@
${\left[\mathbf{M}_{\mathbf{U}}\right]_{ij}\triangleq}\left\{\begin{array}[]{cc}\frac{1}{\gamma_{i}}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{i}\right|^{2},&\mathrm{\ for\quad}i=j,\\ -\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{j}\right|^{2},&\mathrm{\ for\quad}i\neq j.\end{array}\right$. (59)

### v2

@@ III-B Analysis of the Dual Problem
The reformulated problem (30) is convex with strong duality when $\gamma_{k}>0$ and the constraints are strictly feasible. We proceed by taking its Lagrangian dual with respect to (29b). Let $\tilde{\mathbf{B}}_{1},\ldots,\tilde{\mathbf{B}}_{L}\in\mathbb{R}^{(L+1)\times(L+1)}$ denote the dual variables associated with (29b)
@@
which holds because $\mathbf{J}_{\mathbf{R}}$ is positive definite. Thus, the optimal $\mathbf{B}_{\ell}$ must be $\mathbf{B}_{\ell}=\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}^{\mathsf{T}}_{\ell}$ at the saddle-point. Substituting this relation into (33), and again interchanging maximization and minimization, we arrive at
@@
Consider the optimization problem (12). Suppose that: i) $\gamma_{k}>0$ for all $k$; ii) the constraints (12b)-(12c) are strictly feasible; and iii) Assumption 1 holds. Then, $\mathbf{V}^{*}$ is a global optimum solution of (12) if and only if there exists $\boldsymbol{\beta}^{*}\in\mathbb{R}^{L\times L}$ such that $(\mathbf{V}^{*},\boldsymbol{\beta}^{*})$ is a saddle-point solution of
@@
where $\boldsymbol{\beta}\triangleq[\boldsymbol{\beta}_{1},\ldots,\boldsymbol{\beta}_{L}]\in\mathbb{R}^{L\times L}$ and $\mathcal{V}$ denotes the set of feasible beamformers under constraints (12b)-(12c),
@@
$\mathbf{Q}_{\boldsymbol{\beta}}\triangleq\frac{2T}{\sigma_{\text{s}}^{2}}\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right],$ (40)
@@
The development leading to the theorem shows that (38) is the Lagrangian dual of (12) under the stated assumptions. Thus, $\mathbf{V}^{*}$ is an optimal solution of (12) if and only if there exists $\boldsymbol{\beta}^{*}\in\mathbb{R}^{L\times L}$ such that $(\mathbf{V}^{*},\boldsymbol{\beta}^{*})$ is a saddle-point of (38). To further expand the term involving $\mathbf{J}_{\mathbf{V}}$, we note from (6) that
@@
$=\frac{2T}{\sigma_{\text{s}}^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\sum_{ij}[\boldsymbol{\beta}_{\ell}]_{i}\mathbb{E}_{\boldsymbol{\eta}}\left[({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}[\boldsymbol{\beta}_{\ell}]_{j}\right)$ (43b)
$=\frac{2T}{\sigma_{\text{s}}^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[\sum_{i}[\boldsymbol{\beta}_{\ell}]_{i}({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\sum_{j}[\boldsymbol{\beta}_{\ell}]_{j}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43c)
$=\frac{2T}{\sigma_{\text{s}}^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43d)
@@
Examining the max-min problem (39), by construction the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$ is PSD for any $\boldsymbol{\beta}$. Consequently, the trace term can be interpreted as a sum of beamforming powers measured with respect to the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$. Thus, for a fixed $\boldsymbol{\beta}$, the inner problem in (39) is a power maximization along certain directions of interest. Varying $\boldsymbol{\beta}$ corresponds to varying the directions of interest (as illustrated in Fig. 2). Theorem 1 states that there exists a judicious choice $\boldsymbol{\beta}^{*}$ for which the problem of minimizing the BCRB has the same solution as that of maximizing weighted power in the direction of $\mathbf{Q}_{\boldsymbol{\beta}^{*}}$.
@@ III-C AoA Estimation Example
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}=\frac{2T}{\sigma_{\text{s}}^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (45)
@@ IV-A ISAC as Weighted Downlink Power Minimization
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k$ (49b)
@@
$\mathrm{s.t.\ }$~~~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k.$ (51b)
@@
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k.$ (52b)
@@
The condition that $\mathbf{\tilde{Q}}$ must be PSD is important in the result of [12], because the weight matrix $\mathbf{\tilde{Q}}$ for the downlink power minimization problem becomes a noise covariance matrix in the dual uplink channel. Without such a condition, the noise in the uplink problem does not have a physical meaning.
For the ISAC problem, there is no guarantee that for arbitrary $\lambda$ and $\boldsymbol{\beta}$, we would have $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$. In fact, it is essential to properly deal with the cases where $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$ is not PSD for sensing tasks, because the sensing objective typically requires maximizing power in certain spatial directions, rather than minimizing power as in traditional downlink problems. So, $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$ is typically not PSD for the ISAC problem.
Nevertheless, we still need to restrict the space of $(\lambda,\boldsymbol{\beta})$ under consideration. In this paper, we define the following notion of admissibility, and show that the optimal $(\lambda^{*},\boldsymbol{\beta}^{*})$ is always admissible, and moreover, an uplink-downlink duality relationship can be established for all admissible $(\lambda,\boldsymbol{\beta})$’s.
@@
The pair $(\lambda,\boldsymbol{\beta})$ is said to be admissible for a set of channels $(\mathbf{h}_{1},\cdots,\mathbf{h}_{K})$ under the SINR thresholds $(\gamma_{1},\cdots,\gamma_{K})$ for the $K$ users with noise variance $\sigma_{\text{c}}^{2}$, if whenever a set of $\mathbf{V}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}]$ satisfies
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k$ (53)
@@
To understand the notion of admissibility, observe that any $(\lambda,\boldsymbol{\beta})$ for which $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is PSD is obviously admissible, since (54) is satisfied for any set of beamformers $\mathbf{v}_{k}$ (including the ones that satisfy the SINR constraints (53)). Likewise, $(\lambda,\boldsymbol{\beta})$ for which $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is negative definite cannot be admissible.
When $(\lambda,\boldsymbol{\beta})$ is such that $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is neither positive nor negative definite, the notion of admissibility is more intricate. The admissibility condition (54) suggests that $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ behaves somewhat like a PSD matrix, but it is a much weaker condition for two reasons. First, (54) only needs to hold for the set of $\mathbf{v}_{k}$’s satisfying the SINR constraints (53), and not all $\mathbf{v}_{k}$’s. Second, only the sum power is required to be nonnegative; the individual term $\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}$ can actually be negative.
@@
More importantly, we can show that whenever $(\lambda,\boldsymbol{\beta})$ is admissible, an uplink-downlink duality result can be established for the inner minimization in (51), or equivalently (52), as explored in detail in the next section. Such uplink-downlink duality would have been a direct consequence of prior work [12] if $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$, in which case $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ becomes a noise covariance in the dual uplink channel. The interesting and nontrivial fact is that even if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not PSD (so the uplink noise power may be negative in certain directions), as long as $(\lambda,\boldsymbol{\beta})$ is admissible, a duality relationship can still be established.
@@ IV-B Uplink-Downlink Duality
The following theorem states an uplink-downlink duality relationship for admissible $(\lambda,\boldsymbol{\beta})$ pairs. The dual uplink channel here has to satisfy an additional technical condition related to a certain matrix being an $M$-matrix.
@@
Let $\mathcal{A}$ be the set of admissible $(\lambda,\boldsymbol{\beta})$’s for a $K$-user downlink beamforming problem with channels $(\mathbf{h}_{1},\cdots,\mathbf{h}_{K})$, SINR thresholds $(\gamma_{1},\cdots,\gamma_{K})\in\mathbb{R}_{++}^{K}$, and noise variance $\sigma_{\text{c}}^{2}$
@@
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k$ (55b)
where the SINR constraints are assumed to be strictly feasible. The optimal value of this downlink beamforming optimization problem is bounded below if and only if $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$.
Furthermore, for $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, an optimal solution $\mathbf{V}^{*}_{\lambda,\boldsymbol{\beta}}\triangleq\left[\mathbf{v}^{*}_{1},\ldots,\mathbf{v}^{*}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ to (55) can be obtained as
$\mathbf{v}_{k}^{*}=\sqrt{p_{k}^{*}}\mathbf{u}_{k}^{*},\quad\forall k,$ (56)
where $\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}=\left[p^{*}_{1},\ldots,p^{*}_{K}\right]$ is a set of optimal downlink powers given by
$\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}=\sigma_{\text{c}}^{2}\mathbf{M}_{\mathbf{U}_{\lambda,\boldsymbol{\beta}}^{*}}^{-1}\mathbf{1},$ (57)
and $\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}=\left[\mathbf{u}_{1}^{*},\ldots,\mathbf{u}_{K}^{*}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ is a set of normalized beamformers that can be obtained from any optimal solution to the following uplink problem over the beamformers $\mathbf{u}_{1},\ldots,\mathbf{u}_{K}$ and uplink transmit powers $q_{1},\ldots,q_{K}$:
$\underset{\mathbf{U},\;\mathbf{q}}{\mathrm{minimize}}$~~ $\sigma_{\text{c}}^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (58a)
$\mathrm{subject\ to}$~ $\frac{q_{k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}}{\gamma_{k}}\geq\sum_{i\neq k}q_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}+\mathbf{u}_{k}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}_{k},$ (58b)
@@
$\mathbf{M}_{\mathbf{U}}^{-1}\geq 0.$ (58d)
Here, $\mathbf{M}_{\mathbf{U}}\in\mathbb{R}^{K\times K}$ denotes a matrix defined as
@@
${\left[\mathbf{M}_{\mathbf{U}}\right]_{ij}\triangleq}\left\{\begin{array}[]{cc}\frac{1}{\gamma_{i}}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{i}\right|^{2},&\mathrm{\ for\quad}i=j,\\ -\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{j}\right|^{2},&\mathrm{\ for\quad}i\neq j,\end{array}\right$. (59)